North Carolina's Native-American population is 22,000 greater than New York's. New York's Native-American population is 187,000 less than Oklahoma's. If the total population of all three is 437,000, find each state's Native-American population

Answer by Paul(988)

North Carolina = x
New york = y
Oklahoma = z
x=22000+y ---------->x-y=22000
y=z-187000 --------->y-z=-187000
x+y+z=437000

x+y+z=437000
y-z=-187000
x+2y=250000 (new equation)

x+2y=250000 ------------------>x+2y=250000
x-y=22000 -------------------->2x-2y=44000
3x=294000
x=98000

98000-y=22000
y=76000
76000-z=-187000
z=263000
Hence, the population for North Carolina is 98000, for New york its 76000, and for Oklahoma its 263000